OpenStudy (anonymous):

What is the solution to the proportion? --> 12/15 = 8/x

OpenStudy (littlebird):

Multiply 15 by 8 and divide by 12

OpenStudy (anonymous):

x=10

OpenStudy (whpalmer4):

Do you know about cross-multiplication? \[\frac{a}{b} = \frac{c}{d}\]\[a*d = b*c\]

OpenStudy (whpalmer4):

You know all but \(d\), so \[a*d = b*c\]\[\frac{b*c}{a} = d\]

OpenStudy (cwrw238):